From GlueXWiki
Jump to: navigation, search

February 12, 2015 FDC+CDC meeting


  1. Instructions for Bluejeans meeting connection
  2. FDC meeting ID: 290664653
  3. To join via a Web Browser, go to the page [1] https://bluejeans.com/290664653.


  1. Gas issues
  2. Electronics (Fernando, Cody)
  3. Engineering (Dave)
  4. CDC update Resolution Vs. Drift time 1200A(Mike)
  5. FDC update (Lubomir)
  6. Other


Participants: Naomi (CMU) Eugene, Tim, Dave, Fernando, Mike, Chris, Nick, Simon, Beni, and Lubomir (JLab).

Gas issues

- We had very long discussions what to do with the CDC leakage. Everybody agrees that CDC can't be operated like this and the leak has to be fixed. Different proposals: start removing the HVB cards and check with the sniffer the holes at the cheese plate (Beni, Nick, Fernando), check for leaks at the downstream mylar window with a tube on the sniffer (Dave), take some cosmic data overnight, if the leak is at the cheese plate the straws will still have some gas flow and the data should be OK (Lubomir).

- By the time I'm writing the minutes, Nick removed all the cables at the front and a portion of the HVB cards at the bottom. Thus, Mike and Nick were able to sniff the entire cheese plate without finding anything localized. The place with the maximum beeping is at the bottom, between BCAL and CDC, but the gas is heavier than air so it may come from anywhere. On Friday afternoon Tim and Eugene decided to pull the CDC out of the bore. Target people are preparing the target. On Monday morning it will be craned to the floor and the CDC will be pulled out.

- The CO2 sensor came back from repair. We calibrated it with 60/40 CO2/Ar certified mixture in the EEL126, and then started monitoring the CO2 in the first FDC package. Results are linked: some variations for 20 hours, maybe related to the gas temperature, has to be monitored for longer time.


- Fernando: the higher CDC noise (~60mV compared to ~30mV before) related to the loop on HV cables we had before that acted as an inductance. Fernando tested the new HV filter from CAEN on one cable and it worked great: noise <20mV (with or without loop). This was only a prototype box, now we need 13 more to put on all the cables. Fernando agreed that if we remove the CDC, the HV cables can be re-routed with a loop.

- Cody: testing the f125 with the latest firmware at high rates in block-mode. In mode 8 one can reach 5kHz, in mode 1 up to 20kHz, and in mode 7 up to 50kHz - these limits correspond to 100% occupancy.


- Dave: CDC flow sensor is ready.

CDC update

- Mike showed a plot linked above, with the resolutions (sigma of the residuals) vs drift time. The residuals were obtained from tracking in magnetic field (somehow the track quality is poorer without field) after Simon's alignment. The resolutions are very good, ~130 microns for >100ns. The poor resolution for <100ns can be explained by poor timing resolution.

FDC update

- We looked again at the structures seen in the residuals: difference between the cathode reconstructed and the nominal wire position, discussing once more (now with Tim) the possibility for wire offsets. The plots above show both the mean and the R.M.S. values of the residuals as function of x,y and their projections on x (last plot).

- FDC efficiencies for reconstructing 3D hit point in each cell (package 3 shown): it is required to have hits in the two cathodes and the wire plane, the hits to have timing withing 2 samples cathode-to-cathode, and within 48ns wire-to-cathode, as well as the three coordinates to cross at the same point within +/-5mm. The efficiencies are estimated with tracking without magnetic field and shown for 4 different thresholds: 50,100,150,and 200 fADC units. The plateau is at ~95% efficiency. There are places with poorer efficiency due to bad strip/wire channels. In each package, there's also a spot with poorer efficiency which is always at the wire connector side. This maybe due to the timing constraint without taking into account the signal propagation time.